# OffLight: An Offline Multi-Agent Reinforcement Learning Framework for Traffic Signal Control

### v1

@@ 5.2.1 Performance Comparison with Different Levels of Mixing
5.2.2 Effectiveness of Improved Sampling Strategies
@@
We further investigate the contribution of the two components of OffLight: Importance Sampling (IS) and Prioritized Sampling (PS). Both strategies aim to improve learning from offline datasets by addressing different challenges: IS adjusts for the distributional shift between the behavior policy and the target policy, while RBPS focuses the learning on episodes with higher total rewards.
IS consistently outperforms RBPS in all scenarios (Jinan, Hangzhou, Manhattan) and across all demand levels (low, medium, high). The gap between IS and RBPS widens as the traffic demand increases. This is particularly evident in the high traffic demand scenario, where IS achieves up to a 10% improvement in ATT for CQL and 9% improvement for TD3+BC in Manhattan. Medium and high traffic demand scenarios benefit the most from IS, as seen in Hangzhou and Manhattan, where IS outperforms RBPS by several percentage points, especially as congestion increases.
RBPS improves ATT modestly in all scenarios, but its performance is generally lower than IS. In particular, in medium and high traffic demand scenarios, RBPS provides only incremental improvements (around 4-6% in ATT for both CQL and TD3+BC). RBPS performs more consistently across traffic demand levels, but its gains are notably smaller compared to IS, especially in more complex traffic networks like Manhattan.
IS corrects for the distributional shift between the offline dataset (behavior policy) and the target policy, ensuring that transitions more aligned with the target policy are given higher weight. This correction is particularly crucial in high-demand scenarios, where suboptimal data can significantly distort learning. By focusing on more relevant data, IS ensures the model can make more effective decisions, leading to greater improvements in traffic flow, especially in larger, more complex networks like Manhattan. RBPS, while effective in prioritizing episodes with higher rewards, is more sensitive to noise in mixed-quality datasets, especially when suboptimal policies dominate certain episodes. This can lead to smaller performance gains compared to IS, which dynamically adjusts the importance of transitions based on their relevance to the target policy.
5.3 Discussion
In this study, we proposed OffLight, an offline multi-agent reinforcement learning (MARL) framework tailored for traffic signal control (TSC) using heterogeneous, pre-collected datasets. The experiments illustrate that OffLight not only outperforms existing baseline methods but also remains robust across different traffic scenarios and dataset compositions. Key insights from the results include:
Performance with Mixed-Quality Data: OffLight consistently demonstrates superior performance in scenarios involving heterogeneous behavior policies. This is critical for real-world deployments where datasets often comprise both expert and suboptimal policy data. The Graph-GMVAE was crucial in capturing the latent behavior distributions, enabling more accurate policy learning.
Effectiveness of Sampling Strategies: Both importance sampling (IS) and return-based prioritized sampling (RBPS) played pivotal roles in enhancing learning efficiency. IS proved more impactful overall, particularly in high traffic demand scenarios, by correcting for distributional shifts and focusing on transitions more aligned with the target policy. In contrast, RBPS provided more consistent, albeit smaller, gains by prioritizing high-reward episodes. Together, these strategies enabled OffLight to maximize performance, especially in complex environments like Manhattan.
Scalability and Generalization: OffLight scaled well across different network sizes and traffic conditions, from small networks like Jinan to highly complex ones like Manhattan. This demonstrates the framework’s generalizability, making it suitable for deployment in diverse urban settings.
5.3.1 Limitations:
Training Overhead for Graph-GMVAE: Training the Graph-GMVAE, which models the behavior policies, can be computationally expensive, especially in large-scale networks with numerous intersections. This adds to the overall training time and resource requirements, which could limit its real-time deployment capabilities in certain urban environments.
Dependence on Dataset Quality: OffLight’s performance is highly dependent on the quality and diversity of the offline dataset. If the dataset lacks sufficient expert data or is overly biased toward suboptimal policies, OffLight’s learning process may be negatively impacted.
6 Conclusion
@@
OffLight demonstrates the potential to significantly enhance traffic signal control (TSC) in urban networks through offline multi-agent reinforcement learning (MARL). By addressing the key challenges of heterogeneous behavior policies and distributional shifts, OffLight offers a robust solution capable of learning effective policies from mixed-quality datasets. The framework’s incorporation of Graph-GMVAE for behavior policy modeling, alongside importance sampling (IS) and return-based prioritized sampling (RBPS), leads to substantial improvements in traffic efficiency. These advancements reflect the ability of OffLight to scale effectively across different network sizes and traffic conditions, making it a practical tool for offline reinforcement learning for real-world traffic management.
Appendix A OffLight Implementation Details
A.1 Rationale for Averaging Importance Sampling Weights
In multi-agent settings, computing the Importance Sampling (IS) weight as a product across agents can lead to exponential scaling, causing numerical instability and disproportionately amplifying certain transitions. Specifically, the original formulation is given by:
$w_{\text{IS},t}^{i}$ $=\frac{\pi_{\theta}^{i}(A_{t}\mid S_{t})}{\pi_{b}^{i}(A_{t}\mid S_{t})}$ (10)
$=\prod_{i=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid s_{t})}$ (11)
$\approx\prod_{i=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (12)
To mitigate these issues, we adopt an averaging approach:
$w_{\text{IS},t}^{i}$ $\approx\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (13)
Averaging ensures that the combined IS weight remains within a manageable range, preventing the exponential growth associated with multiplication. This approach maintains numerical stability and ensures that each agent’s policy discrepancy contributes proportionally to the overall weight, thereby promoting balanced and stable training dynamics. Empirical results confirm that averaging leads to more stable convergence and prevents any single agent from disproportionately influencing the learning process.
A.2 Pseudocode for Training Behavior Policies
Algorithm 1 summarizes the training procedure of OffLight, highlighting the key steps where Importance Sampling and Return-Based Prioritized Sampling are applied.
Algorithm 1 OffLight Training Procedure
0:  Dataset $\mathcal{D}=\{\tau_{i}\}_{i=1}^{M}$, Target policy parameters $\theta$, Mixer parameters $\phi$
1:  Train GMM-VGAE on $\mathcal{D}$ to model behavior policies
2:  Compute total returns $G_{i}=\sum_{t=0}^{T}r_{t}$ for each trajectory $\tau$
3:  Normalize returns $G_{i}$ to derive RBPS weights $w_{\text{RBPS}}^{i}$
4:  Store estimated behavior policies $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})$ for each episode $i$ in $\mathcal{D}$
5:  Store RBPS weights $w_{\text{RBPS}}^{i}$ for each episode $i$ in $\mathcal{D}$
6:  Initialize target policy parameters $\theta_{\text{target}}\leftarrow\theta$
7:  for each training iteration do
8:     Sample a minibatch of transitions from $\mathcal{D}$ according to RBPS weights $w_{\text{RBPS}}^{i}$
9:     for each transition $(\mathcal{O}_{t}^{i},\mathcal{A}_{t}^{i},\mathcal{R}_{t},\mathcal{O}_{t+1}^{i})$ in the minibatch do
10:        Retrieve stored behavior policy probabilities $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})$ from $\mathcal{D}$
11:        Compute IS weights $w_{\text{IS},t}^{i}$ using:
12:         $w_{\text{IS},t}^{i}=\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$
13:     end for
14:     Combine weights:
15:      $w_{\text{combined},t}^{i}=w_{\text{IS},t}^{i}\times w_{\text{RBPS}}^{i}$
16:     Normalize $w_{\text{combined},t}^{i}$ across the minibatch:
17:      $\tilde{w}_{\text{combined},t}^{i}=\frac{w_{\text{combined},t}^{i}}{\sum_{j}w_{\text{combined},t}^{j}}$
18:     Clamp $\tilde{w}_{\text{combined},t}^{i}$ to prevent extreme values:
19:      $\tilde{w}_{\text{combined},t}^{i}=\text{clamp}(\tilde{w}_{\text{combined},t}^{i},\text{max}=10.0)$
20:     Compute loss $\mathcal{L}$ using:
21:      $\mathcal{L}=\mathbb{E}_{(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\sim\mathcal{D}}\left[\tilde{w}_{\text{combined},t}^{i}\cdot\ell(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\right]$
22:     Update policy parameters $\theta$ using gradient descent on $\mathcal{L}$
23:     Update target policy parameters $\theta_{\text{target}}$ using Polyak averaging or hard updates
24:  end for
A.3 GMM-VGAE Architecture
A.3.1 Encoder: Capturing Spatial and Temporal Dependencies
The encoder processes trajectory data by embedding local observations and actions through linear layers. GATs capture the spatial dependencies among the $N$ traffic intersections by leveraging the graph structure of the traffic network. Subsequently, RNNs, specifically Long Short-Term Memory (LSTM) networks, encode the temporal sequence of these embeddings to capture the action-observation history over time. This dual encoding ensures that both spatial relationships and temporal dynamics are effectively represented in the encoded features.
Formally, the encoding process is defined as: {dmath} μ_t, logσ_t^2 = Encoder(O_t, A_t-1, A), z_t = μ_t + σ_tϵ,  ϵN(0, I), where:
  • $\mathcal{O}$ represents the observations.
  • $\mathcal{A}$ denotes the actions.
  • $\text{Encoder}(\cdot)$ maps the input observations and actions to the parameters of the posterior distribution over the latent variables, yielding the mean $\boldsymbol{\mu}_{t}$ and log-variance $\log\boldsymbol{\sigma_{t}}^{2}$.
  • $\mathbf{z}_{t}$ is the sampled latent variable obtained via the reparameterization trick.
A.3.2 Latent Space: Modeling Behavior Policies
The latent space is structured to represent the distribution of underlying behavior policies using a Gaussian Mixture Model (GMM). By modeling the latent variables $\mathbf{z}$ as a mixture of $K$ Gaussian components, OffLight captures the multimodal nature of behavior policies present in the heterogeneous dataset. This mixture model facilitates the disentanglement of different policy behaviors, enabling accurate estimation and differentiation of effective and suboptimal policies within the data.
The GMM is defined as:
$p(\mathbf{z})$ $=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{z}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}),$ (14)
$q(\mathbf{z}\mid\tau_{i})$ $=\sum_{k=1}^{K}\gamma_{ik}\mathcal{N}(\mathbf{z}\mid\boldsymbol{\mu}_{ik},\boldsymbol{\Sigma}_{ik}),$ (15)
where:
  • $K$ is the number of mixture components.
  • $\pi_{k}$ are the mixture weights with $\sum_{k=1}^{K}\pi_{k}=1$.
  • $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ are the mean and covariance of the $k$-th Gaussian component.
  • $\gamma_{ik}$ represents the responsibility of the $k$-th component for trajectory $\tau_{i}$.
Figure 10: Evaluation of different latent dimensions for modeling behavior policy distribution on Hangzhou scenario
Number of Latent Variables:
The number of latent variables in the Graph-GMM-VGAE model was determined through a validation study using the Hangzhou traffic scenario. We tested several configurations, ranging from 2 to 32 latent dimensions, and measured the reconstruction accuracy of the behavior policies on the validation set. As the number of latent variables increased, the model’s ability to capture heterogeneity in the policy distributions improved. However, beyond 8 latent dimensions, the validation accuracy plateaued, indicating diminishing returns in performance improvement. Additionally, higher-dimensional models risk overfitting, particularly with smaller datasets. Therefore, we selected 8 latent variables as the optimal configuration, offering a good trade-off between accuracy and model complexity.
A.3.3 Decoder: Reconstructing Policy Distributions
The decoder reconstructs the action distributions from the latent variables by concatenating the sampled latent vectors with the encoded observations and passing them through additional GAT layers. This process effectively decodes the policy distributions corresponding to different latent behaviors. The reconstruction ensures that the latent representations accurately capture the decision-making processes of the behavior policies, enabling reliable policy reconstruction and subsequent importance sampling.
The decoding process is formalized as:
$\hat{\mathcal{A}_{t}}$ $=\text{Decoder}(\mathbf{z}_{t},\mathcal{O},\mathcal{A_{t-1}},\mathbf{A}),$ (16)
$p(\mathcal{A}_{t}\mid\mathbf{z}_{t},\mathcal{O}_{t})$ $=\prod_{i=1}^{N}p(a_{t}^{i}\mid\mathbf{z},\mathcal{O}_{t}),$ (17)
where:
  • $\hat{\mathcal{A}}$ represents the reconstructed actions.
  • $\text{Decoder}(\cdot)$ maps the latent variable $\mathbf{z}$ and observations $\mathcal{O}$ back to the action space.
  • $p(a_{t}^{i}\mid\mathbf{z}_{t},\mathcal{O}_{t})$ is modeled as a categorical distribution over possible actions for agent $n$.
OffLight employs a self-supervised learning approach, eliminating the need for explicit labels regarding behavior policies. By minimizing the reconstruction loss alongside the Kullback-Leibler (KL) divergence between the posterior and prior distributions in the latent space, the model learns to encode and decode behavior policies effectively. This self-supervised mechanism ensures that the latent space is structured to accurately reflect the diversity of behavior policies in the dataset.
A.4 Offline RL Framework Details
A.5 Hyperparameters
Component Parameter Value
Encoder LSTM hidden size 128
FC hidden size 128
GAT Layers 3
Attention Heads 4
Latent variable dimensions 8
Hidden layer activation function LeakyReLU
Decoder GAT Layers 3
Attention Heads 4
FC hidden size 128
Hidden layer activation function LeakyReLU
Final layer activation function Softmax
Training Epochs 100
Learning rate 3e-3
Optimizer Adam
Batch size 128
KL divergence weight 1e-4
Table 4: GMM-VGAE Hyperparameters and Architecture Details
Component Parameter Value
Agent Network RNN hidden size 128
FC hidden size 128
GAT Layers 1
Attention Heads 4
Hidden layer activation function ReLU
CQL $\alpha$ 1.0
Target network update interval 200 (episodes)
TD3+BC $\alpha$ 1.0
Target network update $\tau$ 0.005
Training Optimizer Adam
Learning rate 0.001
Table 5: Offline RL Hyperparameters and Architecture Details
We used PyTSC library [28] and the CityFlow simulator [29] to generate traffic scenarios for our experiments, while the OffLight algorithm was implemented using the OffPyMARL framework [30]. Each offline reinforcement learning algorithm was trained for 100 epochs on a system equipped with an Intel(R) Core(TM) i9-9900K CPU @ 3.60GHz and an NVIDIA GeForce RTX 2080 GPU. The entire training process for OffLight took approximately 4 hours per run, with the GMM-VGAE component requiring between 5 to 20 minutes to complete its training.
Appendix B Theoretical Analysis
We analyze the properties of the importance sampling integration. This analysis underscores the robustness and effectiveness of OffLight in offline multi-agent reinforcement learning (MARL) settings, particularly for traffic signal control (TSC) applications.
B.1 Importance Sampling Weight Consistency
In OffLight, importance sampling is integrated to correct for the distributional shift between the behavior policy $\pi_{b}$ and the target policy $\pi_{\theta}$. We aim to show that the importance sampling weights computed using the estimated behavior policy $\hat{\pi}_{b}$ provide consistent and unbiased estimates of expectations under the target policy. For any function $f(\tau)$ of a trajectory $\tau$, the expectation under $\pi_{\theta}$ can be estimated using samples from $\pi_{b}$ as:
@@ B.1 Importance Sampling Weight Consistency
This demonstrates that OffLight’s importance sampling weights yield consistent estimates when the behavior policy is accurately estimated.
B.2 Variance Reduction via Accurate Behavior Policy Estimation
While importance sampling provides unbiased estimates, the variance of the estimator can be high, particularly when the target and behavior policies differ significantly. By accurately estimating $\pi_{b}$ using the GMM-VGAE, OffLight reduces the variance of the importance weights.
Variance Analysis: The variance of the importance sampling estimator is given by:
@@ B.2 Variance Reduction via Accurate Behavior Policy Estimation
Theorem: If $\hat{\pi}_{b}$ is a consistent estimator of $\pi_{b}$, then as the amount of data increases, the variance of the importance sampling estimator decreases.
@@
B.3 Computational Complexity and Scalability
OffLight leverages Graph Neural Networks (GNNs) to capture spatial dependencies among agents. The computational complexity per forward pass is:
$\mathcal{O}(L\cdot(N+E)\cdot d),$ (25)
where: - $L$ is the number of GNN layers. - $N$ is the number of nodes (agents). - $E$ is the number of edges (inter-agent connections). - $d$ is the dimensionality of the embeddings. In sparse traffic networks, $E=\mathcal{O}(N)$, so the complexity scales linearly with $N$. This demonstrates that OffLight is scalable to large networks, such as the Manhattan scenario with 196 intersections.

### v3

@@ V-B1 Performance Comparison with Different Levels of Mixing
V-B2 Effectiveness of Improved Sampling Strategies
@@
Figure 9 analyzes the impact of IS and RBPS on OffLight’s performance. IS improves performance more than RBPS across all scenarios and traffic demand levels, with the performance gap widening as congestion increases. In Manhattan’s high-demand setting, IS improves ATT by up to 10% for CQL and 9% for TD3+BC, demonstrating its effectiveness in mitigating distributional shift. RBPS provides smaller, incremental gains (4-6% ATT reduction), especially in medium and high-demand settings. While RBPS prioritizes high-reward episodes, it is more sensitive to noisy data and less effective than IS in complex networks like Manhattan.
By re-weighting transitions based on their relevance to the target policy, IS ensures that OffLight learns from high-quality data, improving traffic flow decisions in congested scenarios. RBPS, while useful, has a smaller impact when suboptimal policies dominate certain episodes. The combined approach enhances OffLight’s ability to learn from offline datasets efficiently, particularly in large-scale networks.
V-C Discussion
This study introduced OffLight, an offline MARL framework for traffic signal control (TSC) using heterogeneous datasets. OffLight consistently outperforms baselines, demonstrating robustness across different traffic conditions and dataset compositions. By leveraging GMM-VGAE for behavior modeling and incorporating IS and RBPS, OffLight effectively handles mixed-quality data, prioritizing high-reward transitions and mitigating distributional shifts.
OffLight scales well across diverse network sizes, from small to large-scale urban environments. However, its reliance on GMM-VGAE introduces computational overhead, which may limit real-time deployment. Additionally, its performance depends on dataset quality, requiring a balanced mix of expert and suboptimal policies for optimal learning.
Beyond TSC, OffLight’s framework is applicable to smart grids [25] and supply chain management [26], where offline MARL can optimize decision-making in networked systems. Its IS-based policy correction and graph-based modeling make it well-suited for improving coordination in distributed environments.
VI Conclusion
@@
OffLight demonstrates the potential to significantly enhance traffic signal control (TSC) in urban networks through offline multi-agent reinforcement learning (MARL). By addressing the key challenges of heterogeneous behavior policies and distributional shifts, OffLight offers a robust solution capable of learning effective policies from mixed-quality datasets. The framework’s incorporation of GMM-VGAE for behavior policy modeling, alongside IS and RBPS, leads to substantial improvements in traffic efficiency. These advancements reflect the ability of OffLight to scale effectively across different network sizes and traffic conditions, making it a practical tool for offline reinforcement learning for real-world traffic management.
APPENDIX
VI-A Importance Sampling Weight Consistency
In OffLight, IS is integrated to correct for the distributional shift between the behavior policy $\pi_{b}$ and the target policy $\pi_{\theta}$. We aim to show that the importance sampling weights computed using the estimated behavior policy $\hat{\pi}_{b}$ provide consistent and unbiased estimates of expectations under the target policy. For any function $f(\tau)$ of a trajectory $\tau$, the expectation under $\pi_{\theta}$ can be estimated using samples from $\pi_{b}$ as:
@@ VI-A Importance Sampling Weight Consistency
This demonstrates that OffLight’s IS weights yield consistent estimates when the behavior policy is accurately estimated.
VI-B Variance Reduction via Accurate Behavior Policy Estimation
While IS provides unbiased estimates, the variance of the estimator can be high, particularly when the target and behavior policies differ significantly. By accurately estimating $\pi_{b}$ using the GMM-VGAE, OffLight reduces the variance of the importance weights. The variance of the IS estimator is given by:
@@ VI-B Variance Reduction via Accurate Behavior Policy Estimation
Theorem: If $\hat{\pi}_{b}$ is a consistent estimator of $\pi_{b}$, then as the amount of data increases, the variance of the IS estimator decreases.